A sheet of paper with dimensions 22.0 cm by 28.6 cm lies on a desk. Calculate the downward force exerted on the paper by the atmosphere.
Answer #1

given

area, A = 22*28.6 cm^2

= 22*28.6*10^-4 m^2

= 0.06292 m^2

we know, atmospheric pressure, Po = 1.013*10^5 pa


we know, Po = F/A

==> F = Po*A

= 1.013*10^5*0.06292

= 6374 N <<<<<<<<<<------------Answer